OpenStudy (anonymous):

Solve the equation (4+sq rt of -5)(4+sq rt of -2)

OpenStudy (bahrom7893):

well since nobody's doing this.. i guess I'll finish this and leave..

OpenStudy (bahrom7893):

16+4sqrt(-5)+4sqrt(-2)+sqrt(10)

OpenStudy (anonymous):

Thank you [=

OpenStudy (bahrom7893):

16 + 4i*sqrt(5)+4i*sqrt(2)+sqrt(10) 16 + 4i*sqrt(5)+4i*sqrt(2)+sqrt(2)*(sqrt(5)

OpenStudy (bahrom7893):

16 + sqrt(5)(4i+4i*sqrt(2)+sqrt(2)) 16 + sqrt(5)(4i+sqrt(2)(4i+1)) I don't think this can be simplified further.

OpenStudy (bahrom7893):

btw that's not an equation.

OpenStudy (arnavguddu):

(4+ sqrt-5)(4+sqrt-2) = 16+4jsqrt2 + 4jsqrt5 +jsqrt5 x jsqrt2 = 16 + 4j(sqrt 2+sqrt5) -sqrt10